OpenStudy (anonymous):

Complete the square of y=.05x^2-x+1

OpenStudy (tyteen4a03):

Hint: Turn \(\frac{1}{2}x^2 - x\) into \(a(x - h) + k\).

OpenStudy (anonymous):

i think i got it x=10+/-4sqrt 5

OpenStudy (anonymous):

\[y=\frac{ 5 }{100 }x ^{2}-x-1=\frac{ 1 }{ 20 }\left( x ^{2}-20x+100-100 \right)+1\] \[y=\frac{ 1 }{ 20 }\left( x-10 \right)^{2}-5+1\] \[y+4=\frac{ 1 }{ 20 }\left( x-10 \right)^{2}\]

OpenStudy (anonymous):

correction write -1 instead of +1 and then make changes.